# Africanus III. pfb-imaging - a flexible radio interferometric imaging suite

### v1

@@ 1 Introduction
This separation simplifies the problem for a number of reasons. Of particular relevance to the current paper series is the fact that the two problems parallelise completely differently and have very different memory consumption patterns, making it much simpler to efficiently distribute the two problems separately. Although much of the discussion that follows is relevant to both the imaging and calibration problems, this work is mainly concerned with estimating the sky brightness distribution in the presence of known (direction independent) Jones matrices (further details regarding the calibration problem are given in the accompanying paper Kenyon (2024)). We propose an approach based on the preconditioned forward-backward (PFB) algorithm developed in Repetti and Wiaux (2021) and detail an implementation based on the technologies and ideologies discussed in the first paper in this series Perkins (2024). Our proposal is motivated by the fact that the most ubiquitous imaging algorithm in radio interferometry, viz. CLEAN (see e.g. Högbom, 1974; Rau and Cornwell, 2011; Offringa et al., 2014), suffers from a number of limitations.
@@
Our goals for the paper are threefold. First and foremost, we develop a flexible and generalisable framework rooted in forward modelling that is suitable for implementing and accelerating imaging algorithms posed as statistical optimisation problems. One of the principal motivations for creating this framework, dubbed pfb-imaging, is to alleviate the burden on developers aiming to test new imaging algorithms on large interferometric data sets. As discussed in §$\S$ 4, the pfb-imaging framework achieves this by deconstructing the main steps in a typical imaging workflow into separate applications which interact with simpler and more efficient data structures. Secondly, we provide an example of using the framework to implement and accelerate a variant of the Sparsity Averaging Re-weighted Analysis (SARA) family of algorithms (see e.g. Carrillo et al., 2012; Abdulaziz et al., 2019; Thouvenin et al., 2023; Wilber et al., 2023). We show how the hyper-parameters of the algorithm can be coaxed into an intuitive form and provide some guidelines for setting them when dealing with real data. Lastly, we show how our imaging framework ties in with other technologies in the Africanus ecosystem Perkins (2024); Kenyon (2024); Smirnov (2024) and evaluate some of its scaling characteristics in the distributed setting.
@@ 4.1 Overview
In keeping with the philosophy of the current paper series Perkins (2024); Kenyon (2024); Smirnov (2024), pfb-imaging attempts to leverage technologies that are maintained by a large and active open source community as far as possible. Since these technologies are described in detail in the accompanying papers of this series, they are only briefly discussed here.
@@
The previous two papers in the series Perkins (2024); Kenyon (2024) illustrate the scaling of the Dask distributed scheduler using its collections interface to solve embarrassingly parallel problems. It is shown that the scheduler can be coerced (e.g. by carefully cloning root nodes and/or using scheduler plugins) to scale linearly with the amount of resources allocated as long as the problem can be partitioned into a sufficient number of (large enough) chunks. However, the need for regularisation when solving ill-posed inverse problems can make it impossible to partition algorithms in an embarrassingly parallel way and, in general, different forms of regularisation will require different communication patterns between workers. We have observed that the distributed scheduler often fails to respect data locality when attempting to distribute such problems using the collections interface, especially when attempting to embed multiple fan-reduce type operations within a graph. Our attempts to wrangle the collections interface into a form that does not result in unnecessary data transfers between nodes also resulted in unnatural looking code, often obscuring the underlying algorithm.
@@
pfb-imaging interfaces with interferometric data using Dask-MS Perkins (2024). This makes it agnostic to data sources in the form of CASA THE CASA TEAM et al. (2022) Table Data System (CTDS) (van Diepen, 2015), Zarr or Parquet [^11] https://parquet.apache.org/ Apache Software Foundation. (2019) backed measurement sets. The latter two formats allow for efficient parallel read and write operations and are natively compatible with object storage. This opens the door to commodity computing platforms such as AWS while retaining the ability to directly process legacy CTDS backed measurement sets. Images can optionally be rendered to FITS Wells et al. (1981) but special storage has to be provisioned when using computing facilities that rely on object storage since FITS is not natively compatible. The metadata required to produce FITS headers are stored as attributes of the XArray dataset and a dedicated application is available to convert images to FITS files.
In an attempt to keep the implementation as flexible as possible, pfb-imaging comprises a number of distinct applications, each with their own command line interface (CLI). CLIs are constructed using the Click [^12] https://click.palletsprojects.com/ package by using the clickify_parameters decorator available in Stimela2 (see §$\S$ 5.5 of Smirnov (2024)). Applications are grouped under a command group called pfb which is added as an executable during installation. Each application can be run individually from the terminal or grouped together into a highly customisable imager with Stimela2. The cab definitions (see §$\S$ 4 in Smirnov (2024)) for all applications are defined in pfb-imaging, obviating the need for users to provide them manually. Stimela2 can also be used to distribute the imager over a computing cluster. When running in a Kubernetes Brewer (2015) enabled distributed environment, Stimela2 takes care of launching the requested Dask clusters and requires minimal expertise on the part of the user. Importantly, different clusters can be requested for each step in the recipe allowing for efficient use of resources even when running computationally heterogeneous recipes. This is demonstrated using a LocalCluster in the accompanying Stimela2 recipe in B.2 used for some of the results presented in §$\S$ 5. The amendment to this recipe which enables deploying the same workflow on AWS instances is given in Appendix B of Smirnov (2024). Interested readers are encouraged to consult Smirnov (2024) for further details.
@@ 4.2.1 Initialising Stokes visibilities
The init application is responsible for casting data into a form more conducive to the imaging problem. In particular, given the raw data and corresponding net gain table produced by QuartiCal, or corrected data, the individual Stokes visibilities are computed according to equation (8) and the corresponding weights are obtained from the diagonal of (9). Metadata that might be required by other steps in the pipeline (e.g. antenna labels, channel widths and integration time interval) can often be discarded during imaging making the data much more amenable to averaging. Both simple averaging in frequency and baseline dependent averaging in time [^14] Baseline dependent averaging along the frequency axis is not currently supported as it results in ragged arrays which are not currently supported by the wgridder in ducc. are supported and make use of the averaging modules in codex africanus [^15] https://github.com/ratt-ru/codex-africanus Perkins (2024). The degree of averaging is controlled by specifying the number of visibility channels to average together and the level of decorrelation that can be tolerated at some maximum radial distance from the field center (via the max-field-of-view parameter).
@@ 5.1 Initial processing
We validate our imaging framework on observations of the ESO137 field collected using the MeerKAT telescope (project ID SCI-20190418-SM-01). The data consists of two separate observations both spanning approximately 8 hours with 8 seconds integration time and with a spectral resolution of approximately 209 kHz in L-band (0.856 GHz-1.712 GHz) resulting in 4096 frequency channels. The primary calibrator PKS B 1934-638 is sufficiently close to the target field to serve as both bandpass and gain calibrator. Both observations consist of 15 scans where scans alternate between 6 minute observations of PKS B 1934-638 and 1 hour observations of ESO137. We used a combination of auto-flagging with Tricolour [^18] https://github.com/ratt-ru/tricolour and tedious manual inspection to flag the data. The first and last 128 frequency channels were flagged to avoid spectral roll-off at the edges of the band. Frequency channels that are known to be affected by persistent RFI were also flagged. QuartiCal Kenyon (2024) was used to produce initial gain estimates for the target field using the model (39) with the strategy detailed in A. The init application was used to compute Stokes visibilities and weights averaged by a factor of 4 in frequency and using BDA with a maximum decorrelation factor of 0.98 at a field of view of $2^{\circ}$. The resulting dataset is approximately 100GB on disk which amounts to about a factor of 20 in compression. The same level of averaging is used throughout.
@@ 5.3 A reproducible imaging workflow
The main artefacts noticeable in the residuals are the ring-like feature associated with the central source (ESO137-006) and the deep negative holes around certain sources. Closer inspection of the residual image cube revealed that there are marked differences between how well the data are fit between imaging bands, with the worse residuals corresponding to bands that are more affected by RFI. We also found that up-sampling the model in frequency and computing the residuals with a higher frequency resolution eliminates the ring-like feature. This suggests that some of the remaining systematics could be due to the presence of unflagged RFI and the discretisation of the measurement operator (12) along the frequency axis. To test this hypothesis, we selected out the part of the band between 1.295 GHz and 1.503 GHz that is largely free from RFI and imaged it with double the frequency resolution (i.e. 96 channels per band). We also increased the robust weighting factor from -0.3 to 0.5 to suppress short baselines less. Note that the same Stimela2 recipe could be reused simply by overwriting the default parameters from the command line. We deployed the workflow on AWS by augmenting the command line with the additional configuration file that is presented in Appendix B of Smirnov (2024), to which a detailed explanation of how this mechanism works is deferred. Some of the key points to take note of at this stage are the following:
@@
The above functionality can be useful in a number of ways. Firstly, it enables dedicated, observatory specific data centers to offload compute-intensive tasks to commodity computing platforms which are typically equipped with the most up to date hardware. This means that an observatory can plan its data center infrastructure with functionality to service the majority of expected scientific proposals without budgeting for rare and possibly unforeseeable science targets that require additional computing resources. Secondly, it enables developers to test new imaging algorithms without having to copy the full raw data while preserving a perfect record of the transformations the data have undergone thus far. The data selection mechanism also makes it possible to do this on subsets of the data. Finally, by publishing results with an accompanying versioned Stimela2 recipe that is guaranteed to run on a publicly accessible computing platform, it also aids reproducibility. Some of these topics are discussed further in Smirnov (2024).
@@ Acknowledgements
The MeerKAT data used for the benchmarks in this paper series is publicly available via the SARAO archive [^23] https://archive.sarao.ac.za.
@@ Appendix A Calibration
where the complex valued term $B_{p}(\nu)$ is often referred to as the band-pass as it models the frequency response of the antenna, $G_{p}(t)$ is a complex time dependent gain that captures residual time variability in the electronic receiver system and $K_{p}(t,\nu)=\exp(-\imath d(t)(\nu-\nu_{c}))$, with $\nu_{c}$ the central frequency in the band and where $d(t)$ is referred to as the delay, is a phase only term specifically accounting for the fact that there might be time-dependent variability in the path length that the signal has to traverse in going from the individual receivers to the correlator [^24] This uses the fact that a lag in the time domain corresponds to a complex exponential in the Fourier domain i.e. $f(t-t_{0})\leftrightarrow F(\nu)\exp(-2\pi\imath t_{0}\nu)$ for Fourier pairs $f$ and $F$.. With the calibration model specified, QuartiCal can be used to solve for the calibration parameters. The exact implementation details will not concern us here, they are given in the accompanying paper Kenyon (2024). The important thing to note is that QuartiCal splits the calibration problem into distinct time and frequency chunks that do not necessarily align with how the imaging problem is partitioned. This allows it to compute the calibration solutions in a highly efficient and distributed manner. Gain solution intervals (see e.g. Sob et al., 2021)) are used to limit the number of degrees of freedom for the calibration problem and chunks sizes should ideally be chosen to align with the solution intervals. In this work, we solve for the above three terms with a time interval the length of an entire calibrator scan and use a frequency interval the width of the single channel for the band-pass [^25] For the observation considered in §$\S$ 5 the primary calibrator was sufficiently close to the target for the entire duration of the observation so there is only a single calibrator field.. The frequency interval for $K$ is set to the full bandwidth of the data.
@@ Appendix B Stimela2 recipes
This section provides thecommands and Stimela2 recipes used to generate the results presented in this paper. All of the cab definitions (see §$\S$ 4 of Smirnov (2024)) required by the recipes are provided in the cult-cargo [^26] https://github.com/caracal-pipeline/cult-cargo package. We utilised version 0.1.3 for all experiments.
@@ 97 Multiple of the rms to threshold by
This command invokes the same stimela recipe augmented with the kubeconfig.yaml file (provided in Appendix B of Smirnov (2024)) which defines the required Kubernetes configuration to deploy the workflow on AWS. This file also defines the cabs and steps required to retrieve the component model from S3 and compare it with the one produced locally.

### v2

@@ 1 Introduction
This separation simplifies the problem for a number of reasons. Of particular relevance to the current paper series is the fact that the two problems parallelise completely differently and have very different memory consumption patterns, making it much simpler to efficiently distribute the two problems separately. Although much of the discussion that follows is relevant to both the imaging and calibration problems, this work is mainly concerned with estimating the sky brightness distribution in the presence of known (direction independent) Jones matrices (further details regarding the calibration problem are given in the accompanying paper Kenyon et al. (2024)). We propose an approach based on the preconditioned forward-backward (PFB) algorithm developed in Repetti and Wiaux (2021) and detail an implementation based on the technologies and ideologies discussed in the first paper in this series Perkins et al. (2024). Our proposal is motivated by the fact that the most ubiquitous imaging algorithm in radio interferometry, viz. CLEAN (see e.g. Högbom, 1974; Rau and Cornwell, 2011; Offringa et al., 2014), suffers from a number of limitations.
@@
Our goals for the paper are threefold. First and foremost, we develop a flexible and generalisable framework rooted in forward modelling that is suitable for implementing and accelerating imaging algorithms posed as statistical optimisation problems. One of the principal motivations for creating this framework, dubbed pfb-imaging, is to alleviate the burden on developers aiming to test new imaging algorithms on large interferometric data sets. As discussed in §$\S$ 4, the pfb-imaging framework achieves this by deconstructing the main steps in a typical imaging workflow into separate applications which interact with simpler and more efficient data structures. Secondly, we provide an example of using the framework to implement and accelerate a variant of the Sparsity Averaging Re-weighted Analysis (SARA) family of algorithms (see e.g. Carrillo et al., 2012; Abdulaziz et al., 2019; Thouvenin et al., 2023; Wilber et al., 2023). We show how the hyper-parameters of the algorithm can be coaxed into an intuitive form and provide some guidelines for setting them when dealing with real data. Lastly, we show how our imaging framework ties in with other technologies in the Africanus ecosystem Perkins et al. (2024); Kenyon et al. (2024); Smirnov et al. (2024) and evaluate some of its scaling characteristics in the distributed setting.
@@ 4.1 Overview
In keeping with the philosophy of the current paper series Perkins et al. (2024); Kenyon et al. (2024); Smirnov et al. (2024), pfb-imaging attempts to leverage technologies that are maintained by a large and active open source community as far as possible. Since these technologies are described in detail in the accompanying papers of this series, they are only briefly discussed here.
@@
The previous two papers in the series Perkins et al. (2024); Kenyon et al. (2024) illustrate the scaling of the Dask distributed scheduler using its collections interface to solve embarrassingly parallel problems. It is shown that the scheduler can be coerced (e.g. by carefully cloning root nodes and/or using scheduler plugins) to scale linearly with the amount of resources allocated as long as the problem can be partitioned into a sufficient number of (large enough) chunks. However, the need for regularisation when solving ill-posed inverse problems can make it impossible to partition algorithms in an embarrassingly parallel way and, in general, different forms of regularisation will require different communication patterns between workers. We have observed that the distributed scheduler often fails to respect data locality when attempting to distribute such problems using the collections interface, especially when attempting to embed multiple fan-reduce type operations within a graph. Our attempts to wrangle the collections interface into a form that does not result in unnecessary data transfers between nodes also resulted in unnatural looking code, often obscuring the underlying algorithm.
@@
pfb-imaging interfaces with interferometric data using Dask-MS Perkins et al. (2024). This makes it agnostic to data sources in the form of CASA THE CASA TEAM et al. (2022) Table Data System (CTDS) (van Diepen, 2015), Zarr or Parquet [^11] https://parquet.apache.org/ Apache Software Foundation. (2019) backed measurement sets. The latter two formats allow for efficient parallel read and write operations and are natively compatible with object storage. This opens the door to commodity computing platforms such as AWS while retaining the ability to directly process legacy CTDS backed measurement sets. Images can optionally be rendered to FITS Wells et al. (1981) but special storage has to be provisioned when using computing facilities that rely on object storage since FITS is not natively compatible. The metadata required to produce FITS headers are stored as attributes of the XArray dataset and a dedicated application is available to convert images to FITS files.
In an attempt to keep the implementation as flexible as possible, pfb-imaging comprises a number of distinct applications, each with their own command line interface (CLI). CLIs are constructed using the Click [^12] https://click.palletsprojects.com/ package by using the clickify_parameters decorator available in Stimela2 (see §$\S$ 5.5 of Smirnov et al. (2024)). Applications are grouped under a command group called pfb which is added as an executable during installation. Each application can be run individually from the terminal or grouped together into a highly customisable imager with Stimela2. The cab definitions (see §$\S$ 4 in Smirnov et al. (2024)) for all applications are defined in pfb-imaging, obviating the need for users to provide them manually. Stimela2 can also be used to distribute the imager over a computing cluster. When running in a Kubernetes Brewer (2015) enabled distributed environment, Stimela2 takes care of launching the requested Dask clusters and requires minimal expertise on the part of the user. Importantly, different clusters can be requested for each step in the recipe allowing for efficient use of resources even when running computationally heterogeneous recipes. This is demonstrated using a LocalCluster in the accompanying Stimela2 recipe in B.2 used for some of the results presented in §$\S$ 5. The amendment to this recipe which enables deploying the same workflow on AWS instances is given in Appendix B of Smirnov et al. (2024). Interested readers are encouraged to consult Smirnov et al. (2024) for further details.
@@ 4.2.1 Initialising Stokes visibilities
The init application is responsible for casting data into a form more conducive to the imaging problem. In particular, given the raw data and corresponding net gain table produced by QuartiCal, or corrected data, the individual Stokes visibilities are computed according to equation (8) and the corresponding weights are obtained from the diagonal of (9). Metadata that might be required by other steps in the pipeline (e.g. antenna labels, channel widths and integration time interval) can often be discarded during imaging making the data much more amenable to averaging. Both simple averaging in frequency and baseline dependent averaging in time [^14] Baseline dependent averaging along the frequency axis is not currently supported as it results in ragged arrays which are not currently supported by the wgridder in ducc. are supported and make use of the averaging modules in codex africanus [^15] https://github.com/ratt-ru/codex-africanus Perkins et al. (2024). The degree of averaging is controlled by specifying the number of visibility channels to average together and the level of decorrelation that can be tolerated at some maximum radial distance from the field center (via the max-field-of-view parameter).
@@ 5.1 Initial processing
We validate our imaging framework on observations of the ESO137 field collected using the MeerKAT telescope (project ID SCI-20190418-SM-01). The data consists of two separate observations both spanning approximately 8 hours with 8 seconds integration time and with a spectral resolution of approximately 209 kHz in L-band (0.856 GHz-1.712 GHz) resulting in 4096 frequency channels. The primary calibrator PKS B 1934-638 is sufficiently close to the target field to serve as both bandpass and gain calibrator. Both observations consist of 15 scans where scans alternate between 6 minute observations of PKS B 1934-638 and 1 hour observations of ESO137. We used a combination of auto-flagging with Tricolour [^18] https://github.com/ratt-ru/tricolour and tedious manual inspection to flag the data. The first and last 128 frequency channels were flagged to avoid spectral roll-off at the edges of the band. Frequency channels that are known to be affected by persistent RFI were also flagged. QuartiCal Kenyon et al. (2024) was used to produce initial gain estimates for the target field using the model (39) with the strategy detailed in A. The init application was used to compute Stokes visibilities and weights averaged by a factor of 4 in frequency and using BDA with a maximum decorrelation factor of 0.98 at a field of view of $2^{\circ}$. The resulting dataset is approximately 100GB on disk which amounts to about a factor of 20 in compression. The same level of averaging is used throughout.
@@ 5.3 A reproducible imaging workflow
The main artefacts noticeable in the residuals are the ring-like feature associated with the central source (ESO137-006) and the deep negative holes around certain sources. Closer inspection of the residual image cube revealed that there are marked differences between how well the data are fit between imaging bands, with the worse residuals corresponding to bands that are more affected by RFI. We also found that up-sampling the model in frequency and computing the residuals with a higher frequency resolution eliminates the ring-like feature. This suggests that some of the remaining systematics could be due to the presence of unflagged RFI and the discretisation of the measurement operator (12) along the frequency axis. To test this hypothesis, we selected out the part of the band between 1.295 GHz and 1.503 GHz that is largely free from RFI and imaged it with double the frequency resolution (i.e. 96 channels per band). We also increased the robust weighting factor from -0.3 to 0.5 to suppress short baselines less. Note that the same Stimela2 recipe could be reused simply by overwriting the default parameters from the command line. We deployed the workflow on AWS by augmenting the command line with the additional configuration file that is presented in Appendix B of Smirnov et al. (2024), to which a detailed explanation of how this mechanism works is deferred. Some of the key points to take note of at this stage are the following:
@@
The above functionality can be useful in a number of ways. Firstly, it enables dedicated, observatory specific data centers to offload compute-intensive tasks to commodity computing platforms which are typically equipped with the most up to date hardware. This means that an observatory can plan its data center infrastructure with functionality to service the majority of expected scientific proposals without budgeting for rare and possibly unforeseeable science targets that require additional computing resources. Secondly, it enables developers to test new imaging algorithms without having to copy the full raw data while preserving a perfect record of the transformations the data have undergone thus far. The data selection mechanism also makes it possible to do this on subsets of the data. Finally, by publishing results with an accompanying versioned Stimela2 recipe that is guaranteed to run on a publicly accessible computing platform, it also aids reproducibility. Some of these topics are discussed further in Smirnov et al. (2024).
@@ Acknowledgements
The MeerKAT data used for the benchmarks in this paper series is publicly available via the SARAO archive [^23] https://archive.sarao.ac.za. Specifically, the ESO137 and SGRA data are available through the project IDs SCI-20190418-SM-01 and SSV-20200519-FC-01, respectively.
@@ Appendix A Calibration
where the complex valued term $B_{p}(\nu)$ is often referred to as the band-pass as it models the frequency response of the antenna, $G_{p}(t)$ is a complex time dependent gain that captures residual time variability in the electronic receiver system and $K_{p}(t,\nu)=\exp(-\imath d(t)(\nu-\nu_{c}))$, with $\nu_{c}$ the central frequency in the band and where $d(t)$ is referred to as the delay, is a phase only term specifically accounting for the fact that there might be time-dependent variability in the path length that the signal has to traverse in going from the individual receivers to the correlator [^24] This uses the fact that a lag in the time domain corresponds to a complex exponential in the Fourier domain i.e. $f(t-t_{0})\leftrightarrow F(\nu)\exp(-2\pi\imath t_{0}\nu)$ for Fourier pairs $f$ and $F$.. With the calibration model specified, QuartiCal can be used to solve for the calibration parameters. The exact implementation details will not concern us here, they are given in the accompanying paper Kenyon et al. (2024). The important thing to note is that QuartiCal splits the calibration problem into distinct time and frequency chunks that do not necessarily align with how the imaging problem is partitioned. This allows it to compute the calibration solutions in a highly efficient and distributed manner. Gain solution intervals (see e.g. Sob et al., 2021)) are used to limit the number of degrees of freedom for the calibration problem and chunks sizes should ideally be chosen to align with the solution intervals. In this work, we solve for the above three terms with a time interval the length of an entire calibrator scan and use a frequency interval the width of the single channel for the band-pass [^25] For the observation considered in §$\S$ 5 the primary calibrator was sufficiently close to the target for the entire duration of the observation so there is only a single calibrator field.. The frequency interval for $K$ is set to the full bandwidth of the data.
@@ Appendix B Stimela2 recipes
This section provides the commands and Stimela2 recipes used to generate the results presented in this paper. All of the cab definitions (see §$\S$ 4 of Smirnov et al. (2024)) required by the recipes are provided in the cult-cargo [^26] https://github.com/caracal-pipeline/cult-cargo package. We utilised version 0.1.3 for all experiments.
@@ 97 Multiple of the rms to threshold by
This command invokes the same stimela recipe augmented with the kubeconfig.yaml file (provided in Appendix B of Smirnov et al. (2024)) which defines the required Kubernetes configuration to deploy the workflow on AWS. This file also defines the cabs and steps required to retrieve the component model from S3 and compare it with the one produced locally.